OpenStudy (liizzyliizz):

@satellite73 would u be so kinda of to help me realize where I went wrong with this... "solve the differential equation dy/dx- 1/2y cos(x) given that f(0)- 2

OpenStudy (liizzyliizz):

f(0)=2 * and kind * so sorry for the typo

OpenStudy (amistre64):

there is no equation; just a diffy expression.

OpenStudy (liizzyliizz):

I messed up in the beginning I believe when i separated the variables

OpenStudy (liizzyliizz):

y is what screwed me up and *gasp* my homework has Lied to me then -_-

OpenStudy (liizzyliizz):

dy/dx= *

OpenStudy (amistre64):

im assuming its an homogenous thingy \[dy/dx- 1/2y cos(x)=0\] \[dy/dx= 1/2y cos(x)\] \[dy= 1/2y cos(x)dx\] \[2ydy= cos(x)dx\]

OpenStudy (amistre64):

something akin to that i beleive

OpenStudy (amistre64):

so y^2 = sin(x) + C

OpenStudy (amistre64):

when x=0 sin(0) = 0 2^2 = C

OpenStudy (amistre64):

y^2 = sin(x) + 4

OpenStudy (liizzyliizz):

 I had a typo. dy/dx- and dy/dx are two completley different things o.o because I got y^2= sin(x) + 4 when i did it, but i feel like I did something wrong there.